five times the difference of a number and 4 is the same as 4 increased by five times the number plus four times the number. Find the number
Answer by mananth(16949)

let the number bex
5(x-4)= 5x+4x+4
5x-20=9x+4
9x-5x=-20-4
4x=-24
divide by 4
x= -6 the number